OpenStudy (anonymous):

Simplify the expression below. Select one of the options below as your answer: A. 2x^3 + 3x B. 3x^2 - 15 C. 6x^2 + 8x D. 7x^2 - 4x

OpenStudy (anonymous):

12x^3 - 14x^2 - 40x / 2x-5

OpenStudy (anonymous):

thats the problem^

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the thing 2 minutes ago

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ya, and im dumb

OpenStudy (anonymous):

First factor 2x out of the numerator like this:\[\frac{ 12x^3-14x^2-40x }{ 2x-5 }=\frac{ 2x(6x^2-7x-20) }{ 2x-5 }\]Now factor the quadratic in the numerator:\[\frac{ 2x(6x^2-7x-20) }{ 2x-5 }=\frac{ 2x(6x^2-15x+8x-20) }{ 2x-5 }=\frac{ 2x[3x(2x-5)+4(2x-5)] }{ 2x-5 }\]\[=\frac{ 2x(3x+4)\cancel{(2x-5)} }{\cancel{ 2x-5} } =2x(3x+4)=...?\]Do you understand what happened there? We first factored 2x from numerator. Then we factored the remaining quadratic and got 2x-5 in the numerator and denominator so we cancelled those out. We are left with an expression now which can be expanded. Can you expand the remaining expression and tell me what the answer is? @Keandrap15

OpenStudy (mertsj):

Oh for heavens sake!!! Now the genius will take over. Good luck

OpenStudy (anonymous):

lol wat

OpenStudy (anonymous):

none of what u just said makes sense.

OpenStudy (anonymous):

I don't know who you are referring to...*cough cough*

OpenStudy (anonymous):

i dont understand any of what u said in the long paragraph.. i cant learn how to do a problem like this in a paragraph like that @Mertsj was helping just fine

OpenStudy (anonymous):

@Mertsj Just take over lol. I'm sorry for causing any disturbances. He appears to be more comfortable with whatever you were doing =D

OpenStudy (anonymous):

sorry if i sound mean im just frustrated right now and i feels so dumb

OpenStudy (mertsj):

@genius12 Toooooooooooooooooooooooooooooooooooooo Late You wanted it. You got it

OpenStudy (anonymous):

lolololol you meanie @Mertsj

OpenStudy (anonymous):

just forget it then damn

OpenStudy (anonymous):

@Keandrap15 I'll just help you out then. What part of my explanation did you not understand?

OpenStudy (anonymous):

are you there @Keandrap15

OpenStudy (anonymous):

sorry guys im stupid and dont know how to do these kind of problems

OpenStudy (anonymous):

anyone gonna help me please

OpenStudy (anonymous):

nvm i guessed the answer

OpenStudy (anonymous):

got more to answer like this one

OpenStudy (anonymous):

now \[\frac{ (12x ^{3}-14x ^{2}-40x) }{ 2x-5 }\]

OpenStudy (anonymous):

your aim is to factorize above and have answers such that one of the answers can cancel out 2x-5 under

OpenStudy (anonymous):

factorize 2x from above

OpenStudy (anonymous):

\[\frac{ 2x(6x ^{2}-7x-20) }{ 2x-5 }\]

OpenStudy (anonymous):

now we factorize the quadratic equation in bracket above